OpenStudy (anonymous):

PLEASE HELP ASAP A box has 6 beads of the same size, but all are different colors. Tania draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tania would pick a yellow bead on the first draw, then a blue bead, and finally a yellow bead again?

OpenStudy (anonymous):

A box has 6 beads of the same size, but all are different colors. Tania draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tania would pick a yellow bead on the first draw, then a blue bead, and finally a yellow bead again?

OpenStudy (steve816):

\[\frac{ 1 }{ 6 }\times \frac{ 1 }{ 6 }\times \frac{ 1 }{ 6 }=\frac{ 1 }{ 216 }\]

OpenStudy (anonymous):

i dont get this ^

OpenStudy (anonymous):

it dont make sense